Space Stations

NASA Space Station On-Orbit Status 17 October 2017 – Preparing for Friday’s Spacewalk

By Marc Boucher
Status Report
October 19, 2017
Filed under , ,
NASA Space Station On-Orbit Status 17 October 2017 – Preparing for Friday’s Spacewalk
Astronaut Joe Acaba poses between two U.S. spacesuits inside the Quest airlock. He will wear one of those spacesuits on a spacewalk scheduled for Oct. 20, 2017. Joining Acaba on the spacewalk Friday morning will be Commander Randy Bresnik. Credit: NASA.

Two NASA astronauts are getting ready to go on their mission’s third spacewalk on Friday. In the midst of those preparations, the Expedition 53 crew also worked on science gear exploring a wide variety of space phenomena.
Commander Randy Bresnik is preparing to go on the third spacewalk this month with NASA astronaut Joe Acaba. Astronauts Paolo Nespoli and Mark Vande Hei will assist the spacewalking duo in and out of their spacesuits on Friday.

The spacewalkers will replace a camera light on the Canadarm2’s newly-installed Latching End Effector and install a high-definition camera on the starboard truss. Other tasks include the replacement of a fuse on Dextre’s payload platform and the removal of thermal insulation on two electrical spare parts housed on stowage platforms.

Bresnik started his day working on a specialized camera that photograph’s meteors entering the Earth’s atmosphere. Acaba finally wrapped up the day configuring a microscope inside the Fluids Integrated Rack.

Nespoli, from the European Space Agency, set up the new Mini-Exercise Device-2 (MED-2) for a workout session today. Researchers are exploring the MED-2 for its ability to provide effective workouts while maximizing space aboard a spacecraft.

On-Orbit Status Report

Miniature Exercise Device (MED-2): Today crewmembers set up cameras in Node 3 to capture video from multiple views of the Advanced Resistive Exercise Device (ARED) and MED-2 hardware. They also applied body markers, and performed exercises before transferring the video for downlink. The ISS’s exercise equipment is large and bulky, while the MED-2 aims to demonstrate small robotic actuators can provide motion and resistance for crew workout sessions, reducing the size and weight of exercise equipment for long-duration space missions. The MED-2 investigation is a system to test key technologies needed to develop space based exercise equipment that may provide appropriate countermeasures to the adverse effects of microgravity. This technology is critical for the initial design and development of second and third generation Counter Measure Systems (CMS) hardware that is an order of magnitude lighter and smaller than existing ISS class of CMS hardware and that has significantly greater reliability.

Fluids Integrated Rack (FIR) Light Microscopy Module (LMM) Hardware Configuration: The crew continued to conduct the LMM configuration activities that began last week, by removing the Constrained Vapor Bubble (CVB) control box and LMM Monochrome camera, before installing the Fluids and Combustion Facility (FCF) high rate data link, the LMM confocal unit and brackets, and the LMM confocal and wide-field cameras. The LMM hardware configuration will support upcoming ACE-T6 operations. The Light Microscopy Module (LMM) is a modified commercial, highly flexible, state-of-the-art light imaging microscope facility that provides researchers with powerful diagnostic hardware and software onboard the International Space Station (ISS).

Meteor Grating Configuration: The crew removed and replaced gratings in the Meteor camera located in the Window Observational Research Facility (WORF) payload volume. The Meteor payload is a visible spectroscopy instrument with the primary purpose of observing meteors in Earth orbit. Meteor uses image analysis to provide information on the physical and chemical properties of the meteoroid dust, such as size, density, and chemical composition. Since the parent comets or asteroids for most of the meteor showers are identified, the study of the meteoroid dust on orbit provides information about the parent comets and asteroids.

TangoLab-2 Card Troubleshooting: In August, during installation of a payload card into the TangoLab-2 slot, the card did not seat properly. Today as part of the troubleshooting efforts, the crew took a different payload card from TangoLab-1 and then installed it into the slot in TangoLab-2. The crew photographed the inside of TangoLab-2 specifically where connectors are located for downlink.

Veg-03 Operations: The crew photo-documented the status of the plants in the Veggie facility. The Veg-03 investigation uses the Veggie plant growth facility to cultivate a type of cabbage, lettuce and Mizuna which are harvested on-orbit with samples returned to Earth for testing. Organisms grow differently in space, from single-celled bacteria to plants and humans. Future long-duration space missions will require crew members to grow their own food, so understanding how plants respond to microgravity is an important step toward that goal. Veg-03 uses the Veggie plant growth facility to cultivate a type of cabbage, lettuce and Mizuna which are harvested on-orbit with samples returned to Earth for testing.

Earth Imagery from ISS Target Operations: The crew setup and configured the RED camera in the Cupola to capture video footage of New York City and the Nile Delta. The Earth Imagery from ISS investigation creates a series of videos, showcasing Earth from space. These videos will be taken with cameras on the International Space Station in 6K hi-resolution, then integrated into videos for screensavers for public enjoyment, exploration, and engagement.

USOS Extra-Vehicular Activity (EVA) #46 Preparations: Today the crew set up cameras and reviewed robotics procedures/tools in preparation for this Friday’s EVA. The goals of USOS EVA #46 include Removal and Replacement (R&R) of the Latching End Effector (LEE) A Camera Lens Assembly (CLA).

Remote Power Controller Module (RPCM) P12B-B R&R: Today, robotics ground controllers removed a degraded RPCM from external slot P12B-B and attempted to replace it with a functional unit from P13A-G. The installation of the new RPCM into P12B_B slot was unsuccessful this morning. Ground teams have backed off to a park position. Teams are discussing a forward plan. This RPC provides power to the S-band String 2 Transponder.

Today’s Planned Activities
All activities are on schedule unless otherwise noted.

BIOPLENKA. Activation of Cryogem-03 at +37°С setting
KONYUGATSIYA. ТБУ-В No.2 Moding to +29 deg С
Virus Definition File Update on Auxiliary Computer System (ВКС) Laptops
Miniature Exercise Device Operations Session
Onboard Training (OBT) Robotics On-board Trainer (ROBoT) Setup
Regeneration of Micropurification Unit (БМП) Ф2 Cartridge (start)
Review of Robotics procedures for US Extravehicular Activity (EVA) 46
Filling (separation) of ЕДВ [КОВ] for Elektron or ЕДВ-СВ (ЕДВ1 – No.1300 (00071692R, СМ1ПрК), ЕДВ2 – No.1210 (00067598R, ФГБ1ПГО_1_110), ЕДВ3 – No.1242 (00069551R, ФГБ1ПГО_1_111,3.0 CTB for ЕДВ (00053480R))
Meteor On Board Training
METEOR Grating Configuration
PRODUTSENT. Removal of from ТБУ-В No.05 (+4 deg C) and transfer to ТБУ-В No.02 (+29 deg C)
СОЖ Maintenance
Meteor Hard Disk Drive and Antivirus Update
PRODUTSENT. Photography of Bioekologiya case setup in ТБУ-В No.02
KONYUGATSIYA. Removal of Recomb-K from thermostat (+4 deg C) and Insertion into thermostat at +29 deg C. Tagup with specialists
Environmental Health System (EHS) Potable Water Dispenser (PWD) Sample Collect
Crew Medical Officer (CMO) On Board Training (OBT)
BIOPLENKA. Fixation of Constanta No.4-1 cassette biofilm and setup in ТБУ-В No.05. Insertion of Constanta kit for incubation into Cryogem-03. Tagup with specialists
Photo TV GoPro Setup
Bringing ODF up to date using Progress 437 delivered files
KONYUGATSIYA. Process activation, setup in ТБУ-В at +29°С. Tagup with specialists
Water Drainage from SM Rodnik H2O Tank 1 to ЕДВ. SM H2O Tank 1 shell compression
Portable Onboard Computers (POC) Dynamic Onboard Ubiquitous Graphics (DOUG) Setup
On-board Training (OBT) EVA Robotics Onboard Trainer (ROBoT) Session
Inspect ЕДВ cover 1308. Tagup with specialists
IMS Update
KONYUGATSIYA. Completion of activation process, setup in thermostat (+29 deg C). Tagup with specialists
KONYUGATSIYA. Photography of deactivation. Tagup with specialists
VEG-03 Plant Photo
Environmental Health System (EHS) Coliform Water Processing
ENERGY Big Picture Words Reading
TangoLab-2 Card Troubleshooting
ENERGY Equipment check and pre-gather.
CALCIUM. Photo and Placement of Bioecologia Cases. Photo downlink
Preventive maintenance of DC1 АСП hatch sealing mechanisms and Progress 437 [АСА] hatch
Environmental Health System (EHS) Total Organic Carbon Analyzer (TOCA) Potable Water Dispenser (PWD) Sample Analysis
EVA Camera Lens Change
KONYUGATSIYA. Retrieval from thermostat (+29 deg С) and transfer to thermostat (+4 deg С). Tagup with specialists
Verification of ИП-1 Flow Sensor
Payload Locker Empty
Fluids Integrated Rack Rack Doors Open
Take Photo of LAB1O3 Z-Panel to Determine Configuration of all Umbilicals.
LMM Hardware Configuration for Confocal
Exercise Data Downlink via OCA
CONTENT. Experiment Ops
Fill SM Rodnik H2O Water Tank 1 with Disinfectant
Fluids Integrated Rack Rack Doors Close
Fluids Integrated Rack Hardware Return
Glacier Desiccant Swap Alternate
Environmental Health System (EHS) Total Organic Carbon Analyzer (TOCA) Sample Data Record
BIOPLENKA. Removal of Samples from the Thermostat, fixation of Constanta cassette No.4-2 biofilm. Tagup with specialists
BIOPLENKA. Photography during the Insertion into thermostat
Regeneration Micropurification Unit (БМП) Ф2 Absorption Cartridge (end)

Completed Task List Activities
Emergency Mask OBT Questionnaire
Vande Hei Glove Inspection

Ground Activities
All activities are on schedule unless otherwise noted.
Standard commanding

Three-Day Look Ahead:
Wednesday, 10/18: EVA procedure review, Off Duty
Thursday, 10/19: EVA Preparation (tool configuration, procedure conference, cuff checklist preparation, equip lock prep), JEM A/L (MBSU/JOTI removal, Kaber plate install)
Friday, 10/20: USOS EVA #46

QUICK ISS Status – Environmental Control Group:
Component – Status
Elektron – On
Vozdukh – Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) – Off
[СКВ] 2 – SM Air Conditioner System (“SKV2”) – On
Carbon Dioxide Removal Assembly (CDRA) Lab – Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Idle
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Process
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Full Up
Trace Contaminant Control System (TCCS) Node 3 – Off

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.